Tag Archives: FAM

Opticon2 Calibration

Jake and Steven recently noticed localized “hot spots” on most of Jake’s recent qPCRs, where higher levels of fluorescence were consistently showing up in interior portions of the plates than the outer portion of the plates.

Ordered 5nmol of 6-FAM T10 Calibration Standard from Biosearch Technologies and resuspended it in 50μL of 1x dilution buffer (10mM Tris-HCl pH8.0, 50mM NaCl, 5mM MgCl2) to make a 100μM solution. Buffer and dye were stored @ -20C after use.

Buffer calculations: Total Volume = 15mL

  • 1.5mL of 100mM Tris-HCl
  • 150μL of 5M NaCl
  • 750μL of 100mM MgCl2

Made a working dilution of the 6-FAM dye of 300nM in 5mL of 1x dilution buffer (15uL of 100uM dye in 5mL of buffer).

Ran the calibration protocol on the Opticon2 (BioRad) using 50μL of dye in all wells when required by the calibration protocol.

 

Results:

EMPTY PLATE MEASUREMENTS

Empty Plate – Channel 1 voltage measurements

 

Empty Plate – Channel 2 voltage measurements

 

Empty Plate – Ratio of Channel 1 to Channel 2 voltage measurements.

 

The empty plate measurements above show the expected low voltage measurements, but also show a  ~5-fold difference in min/max voltages in each channel. Additionally, the voltage ratios (the third image above) show a wavy pattern, but a smooth, even level from well-to-well is what would be expected if the Opticon was in measuring things properly.

 

DYE PLATE MEASUREMENTS

Dye Measurements – Channel 1 voltage measurements

 

Dye Measurements – Channel 2 voltage measurements

 

Dye Measurements – Channel 1 to Channel 2 voltage measurement ratios.

 

The voltages measured in each channel show the expected increase in voltages relative to the empty plate (> 10x voltage than empty plate). However, the spread between the min/max voltages in both channels is ~4-fold. Additionally, the ratio between the two channels still shows the wavy pattern across all the wells instead of the expected even ratio from well-to-well that should result from the calibration.

It appears the calibration has not resolved the issue.

 

To verify that calibration has failed, I ran two sets of qPCR “protocols” that simply read the dye plate to measure fluorescence across the plate in two plate orientations.

Original Orientation Data File (TAD): 20150630_162622_calibration_test.tad
180 degree rotation Data File (TAD): 20150630_162622_calibration_test_180.tad

 

Dye Fluorescence – Original Orientation

 

Dye Fluorescence – Original Orientation with Fluorescence Graph

 

Dye Fluorescence – 180 Degree Rotation

 

Dye Fluorescence – 180 Degree Rotation with Fluorescence Graph

 

 

First thing to notice is that there’s clearly uneven fluorescence detection across the plate. Viewing the images that also contain the fluorescence graphs reveals a spread of ~8-fold between the highest and lowest fluorescence detection.

The second thing to notice is that, despite rotating the plate 180 degrees, the rotation has no effect on the fluorescence detected in each block location.

Both of these taken together provide strong evidence that there’s an issue with the machine.

Share

Opticon Calibration

Distributed 50uL of FAM calibration dye to wells. Ran out of dye!!

Looking back at old purchasing logs, it turns out we need 2 orders of dye packs to have enough for a 96-well plate.

Will cap existing plate with dye, wrap in foil and store @ -20C.

Ordered an additional pack of dye (Cat# 10006046; not available online, must call BioRad to order). Will ship on Monday. Will finish calibration procedure on Tuesday (20100928). Ugh.

Results:

Empty Plate:

Plate with dye (presumably calibrated):

After running the calibration protocol with the dye, all the wells should show consistent fluorescence levels. Clearly, they do not. Oddly enough, there appears to be a cyclical pattern across the wells of low -> high -> low fluorescence. The calibration protocol advises that if the wells do not exhibit consistent fluorescence across wells, then the plate should be read again. The graph above is the 2nd reading, which appears to be the same as the 1st. Conclusion is that the Opticon 2 is not working correctly and will contact BioRad for price quotes on repairs.

Share